# A 2.9-hour periodic radio transient with an optical counterpart

### v1

@@ Abstract
We present a long-period radio transient (GLEAM-X J 070437) discovered to have an optical counterpart, consistent with a cool main sequence star of spectral type M3. The radio pulsations occur at the longest period yet found, 2.9 hours, and were discovered in archival low-frequency data from the Murchison Widefield Array (MWA). High time resolution observations from MeerKAT show that pulsations from the source display complex microstructure and high linear polarisation, suggesting a pulsar-like emission mechanism occurring due to strong, ordered magnetic fields. The timing residuals, measured over more than a decade, show tentative evidence of a 6 yr modulation. The high Galactic latitude of the system and the M-dwarf star excludes the magnetar interpretation, suggesting a more likely M-dwarf / white dwarf binary scenario for this system.
@@ 1 Introduction
By searching the GaLactic and Extragalactic All-sky Murchison Widefield Array (MWA; Tingay et al., 2013; Wayth et al., 2018) eXtended (GLEAM-X; Hurley-Walker et al., 2022a) survey at 72–231 MHz using a visibility differencing technique, Hurley-Walker et al. (2022b) discovered the radio transient GLEAM-X J1627, which produced 45-Jy, 30–60 s-wide, highly linearly-polarised pulses every 18 minutes for three months in 2018, with a dispersion consistent with a Galactic origin. This was followed by the discovery of GPM J1839, which repeated every 22 minutes, and unlike the previous two very transient sources, was found to be active for at least three decades (Hurley-Walker et al., 2023). Using the Australian SKA Pathfinder (ASKAP) at 888 MHz, Caleb et al. (2024) have discovered a -minute transient ASKAP J1935 which produces both broad, linearly-polarised pulses, and short, circularly-polarised bursts. Dobie et al. (2024) have also detected a single 2-minute pulse from ASKAP J1755, which phenomenologically resembles the other sources, but has not yet been observed to repeat. Dong et al. (2024) have detected a radio source with 421 s, CHIME J0630; at a distance of 170 pc, it is the closest LPT discovered to date, compared to (fairly uncertain) distances of 1–8 kpc for the other sources. de Ruiter et al. (2024) have recently noted linearly-polarised periodic radio emission with a period of 2 hr from a binary M-dwarf / WD system, ILT J11015521.
@@ 2.2 Radio follow-up
with and (at a reference frequency of 1 GHz). Assuming a distance of kpc (Section 2.5), this pulse has a total radio luminosity of erg s-1, where is the beam solid angle at 1 GHz.
The TOAs of individual pulses were determined using the same method as described by Hurley-Walker et al. (2023). The timing residuals (Fig. 1) demonstrate an additional sinusoidal variation of the radio emitter. A timing analysis for 2013–2020 yields a period of s, a period derivative of s s-1, and this additional modulation at days. The spin-down rate should be considered as an upper limit given that its error encapsulates zero spin-down (see bottom panel of Fig. 1).
@@ 2.3 Optical & Infrared counterpart
We performed spectroscopy of this optical counterpart with the Goodman Spectrograph (Clemens et al., 2004) on the SOAR telescope on 2023 October 25 with a 400 line mm-1 grating and a 1.2″-longslit, yielding a FWHM resolution of 6.7 Å over a wavelength range of 4000–7840 Å. Two 30-min exposures were taken, for a total exposure time of 1 hr. The data were reduced, optimally extracted, and combined in the normal manner in IRAF (Tody, 1986).
The SOAR spectrum shows significant molecular lines, such as TiO absorption features (Fig. 3; right), consistent with M stars. Comparing with PHOENIX spectral models (Husser et al., 2013) and X-shooter spectral library DR3 (Verro et al., 2022), we inferred a spectral type of M3 with a temperature of K. Combined with the constraints on the distance (see §2.5) and apparent brightness of the source, which rules out a red giant, we conclude that the optical source is an M3V star. In Fig. 3 (right) we also show the comparison with a WD+M3V system with a WD with K and M⊙ (like AR Sco), showing that a WD star would not be detectable.
@@ 2.4 X-ray observations
During the time XMM-Newton/EPIC was on source, two radio bursts were detected with MeerKAT, at 2023-11-17T02:01:25 and 04:56:29 (UTC). We shifted the X-ray and radio time stamps to the Solar System’s barycenter, using the radio position and the DE-430 JPL ephemeris for reference, and extracted images from various time intervals around these bursts. No excess emission was detected above the background level. We measured upper limits on the EPIC-pn net count rate within a 100-s time window centered on the times of the two bursts, yielding values of 0.09 counts s-1 and 0.06 counts s-1, respectively (3-; 0.3–10 keV). Assuming the same spectral models as those used for the average emission, the corresponding luminosity limits are L erg s-1 and L erg s-1 for the first burst and L erg s-1 and L erg s-1 for the second burst.
@@ 3 Discussion
It is also worth noting that ascribing the absolute peak-to-peak amplitude of 100 s to a change in the compact object’s angular momentum implies a phenomenal transfer of energy, regardless of whether the compact object is a NS or a WD, and they are more likely to arise from changes in the emission region than in glitches of the rotator.
@@ 3.1 Stellar radio flares or exo-planet interaction?
The presence of an M-dwarf coincident with the radio emission is unseen in other long-period radio transients. Radio emission from stars is typically circularly polarized, and appears in flares or bursts due to sunspot activity (Dulk, 1985; Lynch et al., 2017) or, potentially detectable at low radio frequencies (200 MHz), star–exoplanet interactions (Vedantham et al., 2020). Such flares would be wholly undetectable at – pc, and the strong periodicity is also difficult to explain with a purely stellar origin. The closest analogue is CU Virginus, a hot, magnetic chemically-peculiar, Ap-type star 72 pc away, which has been observed to produce pulses of 100 %-circularly-polarised emission at 1–3 GHz, each lasting an hour, and repeating every 12 hours, the rotation rate of the star (Lo et al., 2012). These are thought to be produced from electron cyclotron maser emission in the magnetosphere of the star. The high linear polarisation fraction (up to %) of GLEAM-X J 070437 would be difficult to explain with this emission mechanism. Das & Chandra (2021) observed pulsations from CU Virginus of brightnesses up to 100 mJy at 1–2 GHz; even these are times less luminous than the brightest pulses from GLEAM-X J 070437.
@@ 3.2 A binary system with a neutron star
In the NS case, the high Galactic latitude of the system (), the M-dwarf companion and the faint X-ray emission argue against a magnetar companion. These young NSs are generally isolated, often lie within massive star clusters or supernova remnants, and field decay exponentially dissipates their surface magnetic energy on Myr timescales (Viganò et al., 2013). This time is too short to allow them to leave the plane of the Galaxy even for extreme proper motions. If this source is a NS / M-dwarf binary, it is necessarily an old pulsar with a relatively low field. Many such systems are known in different evolutionary stages, from radio-emitting millisecond pulsars (Lorimer, 2008) to accreting low-mass x-ray binaries (Bahramian & Degenaar, 2023) or transitional systems (Papitto & de Martino, 2022). However in all such cases the pulsar has undergone a recycling period due to accretion from the companion star spinning very fast, with usually hours-timescale orbits. For an NS pulsar binary to have a 2.9 hr spin period while still maintaining enough magnetic field strength or rotational energy to produce radio pulsations, the NS must have been accreting in the propeller regime for an extended period. This process would gradually slow its rotation while preserving its magnetic field, but would require very fine-tuned conditions. Even then, the magnetic field necessary to power the current radio emission (which would have to be spin-down-powered in this case) would exceed 1017 G, which is unrealistically large given the NS old age. Therefore, if this system is indeed a NS binary with a 2.9-hour spin period, it would challenge our current understanding of NS accretion and magnetic field evolution. Alternatively, interpreting the 2.9-hour periodicity as an orbital period introduces more challenges, since the radio transient nature of the system, micro-structures and small duty cycle are difficult to explain through geometric considerations.
@@
Kramer et al. (2024) analysed the quasi-periodicity microstructure timescale of NS pulsations across six orders of magnitude of rotational period, finding a universal relation of ms, with and . This would predict = 15.5 s for GLEAM-X J 070437. Most pulses detected with MeerKAT are about 80 s wide without much distinct substructure visible above the noise. In Fig. 2 we present the brightest detected pulse, and an auto-correlation analysis on data from both the correlator (timescales 2 s–120 s) and PTUSE (1 ms–400 ms). Intermediate timescales were contaminated with instrumental artefacts, likely from the PTUSE gain calibration. On the longer timescales, we see a peak in the ACF at 12 s, broadly consistent with the Kramer et al. prediction.
@@
The detected timescale of periods is much shorter than the Kramer et al. (2024) relation, but can be compared to the spacing of nanoshots observed in the Crab pulsar, periods (cf. Hankins & Eilek, 2007, Fig. 3). However, examples of known nanostructure are rare, and there is no universal relation for nanostructure analogous to the microstructure relation discussed above, so it would be premature to conjecture that the physical cause of the fine structure of GLEAM-X J 070437 is in any way related to that of the Crab nanoshots. The behavior of the fine structure is also comparable to that of PSR J09014046 (Caleb et al., 2022, esp. Extended Data Fig. 2), whose ACFs indicate quasi-periodicities on similar timescales (i.e. tens of milliseconds), even though the ratio of the quasi-periodic timescale to the rotation period is two orders of magnitude larger than for GLEAM-X J 070437.
@@ 3.3 A binary system with a white dwarf
Another possibility is that the compact object in the binary system is a WD, presumably with a strong magnetic field. In polars, a type of magnetic cataclysmic variable (MCV), the magnetic field of the white dwarf is 10–200 MG, and the two stars are locked in synchronous rotation. Intermediate polars have slightly weaker magnetic fields (1–10 MG); the WD rotates at a faster rate than the orbit. With an even faster WD rotation rate, AR Sco-type systems consist of an M-dwarf/WD pair in a tight (-hr) orbit, such that the M-dwarf is nearly Roche-lobe filling (Marsh et al., 2016). Interactions between the two stars cause pulsar-like radio emission from the magnetosphere of the WD, which has been directly observed in J1912-44, as the beam of the WD crosses our line-of-sight (Pelisoli et al., 2023). The two known “WD pulsar” systems were selected by looking for high optical variability, which preferentially selects for nearby systems with quickly-spinning WDs (minutes). It is therefore possible that slowly-rotating AR Sco-type systems could have been missed by optical searches. The observed periodicity of hr is however consistent with the rotation rate of the WDs in the MCV population (Norton et al., 2004).
The recently discovered ILT J11015521 de Ruiter et al. (2024), a hr synchronized polar system composed by a WD and an M-dwarf showing periodic and coherent radio emission, might be a twin system with GLEAM-X J 070437. In this scenario, the 2.9 hr periodicity might be interpreted as both the spin and the orbital period, and the yr modulation either as timing noise or as a possible super-orbital modulation.
The UV limits derived by a 3 ks of Swift observations yielded upper limits on the UV/blue emission, shown in Fig. 3. This rules out the presence of WDs with 13,000 K, i.e. about 44% of the magnetic WD population. However, the WDs in J1912-44, AR Sco and ILT J11015521 are all cooler, with the latter having a temperature of K (de Ruiter et al., 2024). Sensitive observations below a wavelength of 400 nm are required.
@@
In the hypothesis of the long yr periodicity being instead the orbit of the system, the small amplitude of the timing residuals (50s) implies a very low inclination angle (Fig. 4). The beam opening angle of this system would then be extremely small, just 30 s%. If past accretion from the stellar companion has aligned the rotation and orbital axes, as is common in MCVs, then the radio beam would also be nearly aligned with the WD rotation axis. This would then present a selection effect: radio searches will tend to find systems with small inclination angles.
Assuming an M3V stellar mass of 0.32 and a WD mass of 0.8, the Roche lobe radius of the M3V star is 0.2 A.U. For , and a circular day orbit, the distance between the two sources would be 2.4 A.U. So, potentially, rather than an AR Sco type scenario with a close-by interaction, or a polar-like system as ILT J11015521, the pulsar emission is instead generated by a stellar wind flowing from the M-dwarf companion onto the magnetosphere of the WD, where it is accelerated, producing radio emission.

### v2

@@ Abstract
We present a long-period radio transient (GLEAM-X J 070437) discovered to have an optical counterpart, consistent with a cool main sequence star of spectral type M3. The radio periodicity occurs at the longest period yet found, 2.9 hours, and was discovered in archival low-frequency data from the Murchison Widefield Array (MWA). High time resolution observations from MeerKAT show that pulsations from the source display complex microstructure and high linear polarisation, suggesting a pulsar-like emission mechanism occurring due to strong, ordered magnetic fields. The timing residuals, measured over more than a decade, show tentative evidence of a 6 yr modulation. The high Galactic latitude of the system and the M-dwarf star excludes a magnetar interpretation, suggesting a more likely M-dwarf / white dwarf binary scenario for this system.
@@ 1 Introduction
By searching the GaLactic and Extragalactic All-sky Murchison Widefield Array (MWA; Tingay et al., 2013; Wayth et al., 2018) eXtended (GLEAM-X; Hurley-Walker et al., 2022a) survey at 72–231 MHz using a visibility differencing technique, Hurley-Walker et al. (2022b) discovered the radio transient GLEAM-X J1627, which produced 45-Jy, 30–60 s-wide, highly linearly-polarised pulses every 18 minutes for three months in 2018, with a dispersion consistent with a Galactic origin. This was followed by the discovery of GPM J1839, which repeated every 22 minutes, and unlike the previous two very transient sources, was found to be active for at least three decades (Hurley-Walker et al., 2023). Using the Australian SKA Pathfinder (ASKAP) at 888 MHz, Caleb et al. (2024) have discovered a 54-min transient ASKAP J1935 which produces both broad, linearly-polarised pulses, and short, circularly-polarised bursts. Dobie et al. (2024) have also detected a single 2-minute pulse from ASKAP J1755, which phenomenologically resembles the other sources, but has not yet been observed to repeat. Dong et al. (2024) have detected a radio source with 421 s, CHIME J0630; at a distance of 170 pc, it is the closest LPT discovered to date, compared to (fairly uncertain) distances of 1–8 kpc for the other sources. de Ruiter et al. (2024) have recently noted linearly-polarised periodic radio emission with a period of 2 hr from a binary M-dwarf / WD system, ILT J11015521, interpreting it as the spin and orbital periodicity of a polar-like system.
@@ 2.2 Radio follow-up
with and (at a reference frequency of 1 GHz). This pulse has a total radio luminosity of erg s-1, where is the beam solid angle at 1 GHz, and is the source distance (see Section 2.5).
The TOAs of individual pulses were determined using the same method described by Hurley-Walker et al. (2023). The timing residuals (Fig. 1) demonstrate an additional sinusoidal variation of the radio emitter. A timing analysis for 2013–2020 yields a period of s, a period derivative of s s-1, and this additional modulation at days. The spin-down rate should be considered as an upper limit given that its error encapsulates zero spin-down (see bottom panel of Fig. 1).
@@ 2.3 Optical & Infrared counterpart
We performed spectroscopy of this optical counterpart with the Goodman Spectrograph (Clemens et al., 2004) on the SOAR telescope on 2023 October 25 with a 400 line mm-1 grating and a 1.2″-longslit, yielding a FWHM resolution of 6.7 Å over a wavelength range of 4000–7840 Å. Two 30-min exposures were taken, for a total exposure time of 1 hr. The data were reduced, optimally extracted, and combined in the normal manner in IRAF (Tody, 1986) and the spectra were flux-calibrated using a spectrum of the spectrophotometric standard EG 274 via the IRAF tasks standard, sensfunc, and calibrate.
The SOAR spectrum shows significant molecular lines, such as TiO absorption features (Fig. 3; right), consistent with M stars. Comparing with PHOENIX spectral models (Husser et al., 2013) and X-shooter spectral library DR3 (Verro et al., 2022), we inferred a spectral type of M3 with a temperature of K. Combined with the constraints on the distance (see §2.5) and apparent brightness of the source, which rules out a red giant, we conclude that the optical source is an M3V star. In Fig. 3 (right) we also show the comparison with a WD+M3V system with a WD with K and M⊙ (like AR Sco), showing that such a WD star would not be detectable (see Section 3.3 for a more thorough treatment of temperature upper limits).
To search for UV emission from a potential WD, we observed GLEAM-X J 070437 with Swift/UVOT (Gehrels et al., 2004; Roming et al., 2005) for 4ks on 2023-06-15 (Obs. ID. 00016076001) with UVOT in 0x308f mode (equal weighting of UV filters). We reduced and analyzed the data following standard procedures.222https://www.swift.ac.uk/analysis/uvot/index.php We found no evidence for UV emission at the location of the source, with 3- Vega magnitude upper limits of UM2 = 20.4 , UW1 = 20.5 , UW2 = 20.6. These upper limits are plotted in Figure 3.
@@ 2.4 X-ray observations
During the time XMM-Newton/EPIC was on source, two radio bursts were detected with MeerKAT, at 2023-11-17T02:01:25 and 04:56:29 (UTC). We shifted the X-ray and radio time stamps to the Solar System’s barycenter, using the radio position and the DE-430 JPL ephemeris for reference, and extracted images from various time intervals around these bursts. No excess emission was detected above the background level. We measured upper limits on the EPIC-pn net count rate within a 100-s time window centered on the times of the two bursts, yielding values of 0.09 counts s-1 and 0.06 counts s-1, respectively (3-; 0.3–10 keV). Assuming the same spectral models as those used for the average emission, the corresponding luminosity limits are L erg s-1 and L erg s-1 for the first burst and L erg s-1 and L erg s-1 for the second burst. Note that all luminosities are dependent on the distance via , which currently has significant uncertainty (see below).
@@ 3 Discussion
It is also worth noting that ascribing the absolute peak-to-peak amplitude of 100 s to a change in the compact object’s angular momentum implies a phenomenal transfer of energy, regardless of whether the compact object is a NS or a WD. Such a significant energy transfer seems improbable, suggesting that the observed variations more likely arise from changes in the emission region rather than rotational glitches in the compact object itself.
@@ 3.1 Stellar radio flares or exo-planet interaction?
The presence of an M-dwarf coincident with the radio emission is unseen in other long-period radio transients. Radio emission from stars is typically circularly polarized, and appears in flares or bursts due to sunspot activity (Dulk, 1985; Lynch et al., 2017) or, potentially detectable at low radio frequencies (200 MHz), star–exoplanet interactions (Vedantham et al., 2020). Such flares would be wholly undetectable at – pc, and the strong periodicity is also difficult to explain with a purely stellar origin. The closest analogue is CU Virginus, a hot, magnetic chemically-peculiar, Ap-type star 72 pc away, which has been observed to produce pulses of 100 %-circularly-polarised emission at 1–3 GHz, each lasting an hour, and repeating every 12 hours, the rotation rate of the star (Lo et al., 2012). These are thought to be produced from electron cyclotron maser emission in the magnetosphere of the star. The high linear polarisation fraction (up to %) of GLEAM-X J 070437 would be difficult to explain with this emission mechanism. Das & Chandra (2021) observed pulsations from CU Virginus of brightnesses up to 100 mJy at 1–2 GHz; even these are 400 times less luminous than the brightest pulses from GLEAM-X J 070437.
@@ 3.2 A binary system with a neutron star
In the NS case, the high Galactic latitude of the system (), the M-dwarf companion and the faint X-ray emission argue against a magnetar companion. These young NSs are generally isolated, often lie within massive star clusters or supernova remnants, and field decay exponentially dissipates their surface magnetic energy on Myr timescales (Viganò et al., 2013). This time is too short to allow them to leave the plane of the Galaxy even for extreme proper motions. If this source is a NS / M-dwarf binary, then the NS is necessarily an old pulsar with a relatively low field. Many such systems are known in different evolutionary stages, from radio-emitting millisecond pulsars (Lorimer, 2008) to accreting low-mass x-ray binaries (Bahramian & Degenaar, 2023) or transitional systems (Papitto & de Martino, 2022). However in all such cases the pulsar has undergone a recycling period due to accretion from the companion star spinning very fast, with usually hours-timescale orbits. For an NS pulsar binary to have a 2.9 hr spin period while still maintaining enough magnetic field strength or rotational energy to produce radio pulsations, the NS must have been accreting in the propeller regime for an extended period. This process would gradually slow its rotation while preserving its magnetic field, but would require very fine-tuned conditions. Even then, the magnetic field necessary to power the current radio emission would exceed 1017 G, which is unrealistically large for an aged NS required for this scenario. The spin-down luminosity erg s-1 would still be at least an order of magnitude too low to explain the radio luminosity. Therefore, if this system is indeed a NS binary with a 2.9-hour spin period, it would challenge our current understanding of NS accretion and magnetic field evolution. Alternatively, interpreting the 2.9-hour periodicity as an orbital period introduces more challenges, since the radio transient nature of the system, micro-structures and small duty cycle are difficult to explain through geometric considerations.
@@
Kramer et al. (2024) analysed the quasi-periodicity microstructure timescale of NS pulsations across six orders of magnitude of rotational period, finding a universal relation of ms, with and . This would predict = 15.5 s for GLEAM-X J 070437. Most pulses detected with MeerKAT are a maximum of 80 s wide without much distinct substructure visible above the noise. In Fig. 2 we present the brightest detected pulse, and an auto-correlation analysis on data from both the correlator (timescales 2 s–120 s) and PTUSE (1 ms–400 ms). Intermediate timescales were contaminated with instrumental artefacts, likely from the PTUSE gain calibration. On the longer timescales, we see a peak in the ACF at 12 s, broadly consistent with the Kramer et al. prediction.
@@
On shorter timescales, the observed peak at periods is much shorter than the Kramer et al. (2024) relation, but can be compared to the spacing of nanoshots observed in the Crab pulsar, periods (cf. Hankins & Eilek, 2007, Fig. 3). However, examples of known nanostructure are rare, and there is no universal relation for nanostructure analogous to the microstructure relation discussed above, so it would be premature to conjecture that the physical cause of the fine structure of GLEAM-X J 070437 is in any way related to that of the Crab nanoshots. The behavior of the fine structure is also comparable to that of PSR J09014046 (Caleb et al., 2022, esp. Extended Data Fig. 2), whose ACFs indicate quasi-periodicities on similar timescales (i.e. tens of milliseconds), even though the ratio of the quasi-periodic timescale to the rotation period is two orders of magnitude larger than for GLEAM-X J 070437.
@@ 3.3 A binary system with a white dwarf
Another possibility is that the compact object in the binary system is a WD, presumably with a strong magnetic field. The spin-down luminosity in this scenario would be erg s-1, potentially more easily driving the observed radio emission. In polars, a type of magnetic cataclysmic variable (MCV; see Warner, 1995, for a review), the magnetic field of the white dwarf is 10–200 MG, and the two stars are locked in synchronous rotation. Intermediate polars (IPs) have slightly weaker magnetic fields (1–10 MG); the WD rotates at a faster rate than the orbit. With an even faster WD rotation rate, AR Sco-type systems consist of an M-dwarf/WD pair in a tight (-hr) orbit, such that the M-dwarf is nearly Roche-lobe filling (Marsh et al., 2016). Interactions between the two stars cause pulsar-like radio emission from the magnetosphere of the WD, which has been directly observed in J1912-44, as the beam of the WD crosses our line-of-sight (Pelisoli et al., 2023). The two known “WD pulsar” systems were selected by looking for high optical variability, which preferentially selects for nearby systems with quickly-spinning WDs (minutes). It is therefore possible that slowly-rotating AR Sco-type systems could have been missed by optical searches. The observed periodicity of hr might be consistent with both the rotation rate and/or the orbit of the WD in the MCV scenario (Norton et al., 2004).
@@
In the hypothesis of the long yr periodicity being the orbit of the system, the small amplitude of the timing residuals (50s) implies a very low inclination angle (Fig. 4). The beam opening angle of this system would be extremely small, just 30 s%. No MCVs have been ever observed with such long orbital period, although long period binary WD systems exist (Ciardullo et al., 1999).
The recently discovered ILT J11015521 (de Ruiter et al., 2024), a hr synchronized polar system comprising a WD and an M-dwarf showing periodic and coherent radio emission, might be a twin system to GLEAM-X J 070437. In this scenario, the 2.9 hr periodicity might be interpreted as both the spin and the orbital period, and the yr modulation either as timing noise or as a possible super-orbital modulation. The polarisation properties of this system are similar to GLEAM-X J 070437, with a mix of linear and circular polarisation, which could be explained by a relativistic version of the electron cyclotron maser mechanism (Qu & Zhang, 2024).333The polarisation properties of AR Sco published by Buckley et al. (2017) demonstrate the high-frequency radio properties of the interaction between the WD and the M-dwarf, rather than the WD pulsar beam, and are therefore not directly comparable. Pelisoli et al. (2023) do not describe the radio polarisation properties of the WD pulsar beam of J 1912. We note that an orbital period of 2.9 hr lies in the orbital period gap for MCVs (Knigge et al., 2011), something that have been observed for some polars, not for IPs. On the other hand, if interpreted as the spin period of an IP with a longer orbital period, this would represent the slowest IP ever discovered. We therefore tentatively suggest that GLEAM-X J 070437 might be a polar system, similar to ILT J11015521. Assuming an M3V stellar mass of 0.32 and a WD mass of 0.8, the Roche lobe radius of the M3V star is 0.2 A.U. For a 2.9 hr circular orbit with identical masses, the orbital radius would be 0.005 AU, placing the WD well within the Roche lobe.
The UV limits derived by a 4 ks of Swift observations yielded upper limits on the UV/blue emission, shown in Fig. 3. This rules out the presence of a hot WDs with 20,000 K, i.e. about 20% of the magnetic WD population. Furthermore, no strong emission lines are seen in the infrared spectrum, which could be expected for a hot WD in a close orbit. However, the WDs in J1912-44, AR Sco and ILT J11015521 are all much cooler, with the latter having a temperature of K (de Ruiter et al., 2024). The upper limit we derive from the infrared spectrum and the UV observations on the presence of a WD in the system are not strongly constraining. Sensitive observations below a wavelength of 400 nm are required to determine whether a WD is present.